# The theory of implicit operations

### v1

@@ 8. Extendable implicit operations
$\mathsf{ext}_{pp}(\mathsf{M},\mathsf{K})=\mathsf{ext}(\mathsf{M},\mathsf{K})\cap\mathsf{imp}_{pp}(\mathsf{K})\,\,\text{ and }\,\,\mathsf{ext}_{eq}(\mathsf{M},\mathsf{K})=\mathsf{ext}(\mathsf{M},\mathsf{K})\cap\mathsf{imp}_{eq}(\mathsf{K}).$
When $\mathsf{M}=\mathsf{K}$, we write $\mathsf{ext}(\mathsf{K}),\mathsf{ext}_{pp}(\mathsf{K})$, and $\mathsf{ext}_{eq}(\mathsf{K})$ instead of $\mathsf{ext}(\mathsf{K},\mathsf{K}),\mathsf{ext}_{pp}(\mathsf{K},\mathsf{K})$, and $\mathsf{ext}_{eq}(\mathsf{K},\mathsf{K})$. Moreover, when an implicit operation is in $\mathsf{ext}(\mathsf{K})$, we simply say it is extendable.
@@
The proof of 8.4 hinges on the following observation.
@@
Recall from 3.9 that each implicit operation $f$ of $\mathsf{K}$ is defined by an existential positive formula $\varphi_{f}$. Then consider the following set of formulas in the language of $\mathsf{K}$ expanded with fresh constants $\{c_{a}:a\in A\}$ for the elements of $A$:
@@
Now, we proceed to prove the first part of the statement of 8.4. Consider $\bm{A}\in\mathsf{K}$. We will define a sequence $\{\bm{A}_{i}:i\in\mathbb{N}\}$ of members of $\mathsf{K}$. First, let $\bm{A}_{0}=\bm{A}$. Then suppose $\bm{A}_{i}\in\mathsf{K}$ has already been defined. By Claim 8.6 there exists $\bm{A}_{i+1}\in\mathsf{K}$ with $\bm{A}_{i}\leqslant\bm{A}_{i+1}$ and $A_{i}^{n}\subseteq\mathsf{dom}(f^{\bm{A}_{i+1}})$ for each $n$-ary $f\in\mathsf{ext}(\mathsf{K})$. By definition the sequence $\{\bm{A}_{i}:i\in\mathbb{N}\}$ constructed in this way is such that
@@
Now, as $\mathsf{K}$ is a universal class, it is closed under unions of chains of algebras by 2.4. Therefore, the union $\bm{B}$ of the chain in the above display belongs to $\mathsf{K}$. Furthermore, $\bm{A}=\bm{A}_{0}\leqslant\bm{B}$. To conclude the proof of the first part of the statement, it only remains to show that $f^{\bm{B}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. To this end, let $f\in\mathsf{ext}(\mathsf{K})$ be $n$-ary and $b_{1},\dots,b_{n}\in B$. As $\bm{B}$ is the union of the chain in the above display, there exists $i\in\mathbb{N}$ such that $b_{1},\dots,b_{n}\in A_{i}$. By the definition of $\bm{A}_{i+1}$ we have $\langle b_{1},\dots,b_{n}\rangle\in A_{i}^{n}\subseteq\mathsf{dom}(f^{\bm{A}_{i+1}})$. As $\bm{A}_{i+1}\leqslant\bm{B}$, we can apply Proposition 8.1, obtaining $\langle b_{1},\dots,b_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$. Hence, $f^{\bm{B}}$ is a total operation, as desired.
To prove the second part of the statement of 8.4, suppose that $\mathsf{K}$ is a quasivariety and consider $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$. In view of the first part of the statement of 8.4, there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $f^{\bm{B}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. By Proposition 8.5 there also exist $\bm{C}\in\mathsf{K}_{\textsc{rsi}}$ with $\bm{A}\leqslant\bm{C}$ and a surjective homomorphism $h\colon\bm{B}\to\bm{C}$. To conclude the proof, it only remains to show that $f^{\bm{C}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. To this end, consider an $n$-ary $f\in\mathsf{ext}(\mathsf{K})$ and $c_{1},\dots,c_{n}\in C$. Since $h\colon\bm{B}\to\bm{C}$ is surjective, there exist $b_{1},\dots,b_{n}\in B$ such that $h(b_{j})=c_{j}$ for each $j\leqslant n$. Recall that $f^{\bm{B}}$ is total because $f\in\mathsf{ext}(\mathsf{K}^{\textup{fg}},\mathsf{K})$. Therefore, $\langle b_{1},\dots,b_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$. As $f$ is an implicit operation and $h$ a homomorphism, we conclude that $\langle c_{1},\dots,c_{n}\rangle=\langle h(b_{1}),\dots,h(b_{n})\rangle\in\mathsf{dom}(f^{\bm{C}})$. ∎
The following is a consequence of 8.4.
@@
Given a universal class $\mathsf{K}$, the classes $\mathsf{ext}(\mathsf{K})$ and $\mathsf{ext}_{pp}(\mathsf{K})$ are closed under composition.
@@
Consider an $n$-ary $g\in\mathsf{ext}(\mathsf{K})$ and $m$-ary $f_{1},\dots,f_{n}\in\mathsf{ext}(\mathsf{K})$. Let $\bm{A}\in\mathsf{K}$. By 8.4 there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $g^{\bm{B}},f_{1}^{\bm{B}},\dots,f_{n}^{\bm{B}}$ are total. It then follows from the definition of composition that $g(f_{1},\dots,_{n})^{\bm{B}}$ is also total. Therefore, $\mathsf{ext}(\mathsf{K})$ is closed under composition. As $\mathsf{imp}_{pp}(\mathsf{K})$ is closed under composition by 3.13, we obtain that $\mathsf{ext}_{pp}(\mathsf{K})$ is also closed under composition because $\mathsf{ext}_{pp}(\mathsf{K})=\mathsf{ext}(\mathsf{K})\cap\mathsf{imp}_{pp}(\mathsf{K})$. ∎
@@
Recall from 4.16 that the class of cancellative commutative monoids forms a quasivariety, which we denote by $\mathsf{CCMon}$. The importance of cancellative commutative monoids is due to the following well-known result (see, e.g., [84, pp. 39–40]).
@@
    if $\mathbb{P}(\mathsf{K})\subseteq\mathsf{K}\subseteq\mathbb{Q}(\mathsf{M})$, then $\mathsf{ext}_{pp}(\mathsf{K})=\mathsf{ext}_{pp}(\mathsf{M},\mathsf{K})$.
@@
(i): Suppose that $\mathsf{K}\subseteq\mathbb{U}(\mathsf{M})$. The inclusion from left to right follows from 8.3. To prove the other inclusion consider $f\in\mathsf{ext}(\mathsf{M},\mathsf{K})$, $\bm{A}\in\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. By 2.2 we have $\mathbb{U}(\mathsf{M})=\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Together with $\bm{A}\in\mathsf{K}\subseteq\mathbb{U}(\mathsf{M})$, this yields $\bm{A}\in\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Therefore, there exist $\bm{B}\in\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$ and an embedding $h\colon\bm{A}\to\bm{B}$. By Claim 8.12 there exists also $\bm{C}\in\mathsf{K}$ such that $\bm{B}\leqslant\bm{C}$ and $\langle h(a_{1}),\dots,h(a_{n})\rangle\in\mathsf{dom}(f^{\bm{C}})$.
@@
(ii): Suppose that $\mathbb{P}(\mathsf{K})\subseteq\mathsf{K}\subseteq\mathbb{Q}(\mathsf{M})$. The inclusion from left to right follows from 8.3. To prove the other inclusion consider $f\in\mathsf{ext}_{pp}(\mathsf{M},\mathsf{K})$, $\bm{A}\in\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. Assume that $f$ is defined by a pp formula $\varphi$. By Theorem 2.2 we have $\mathbb{Q}(\mathsf{M})=\mathbb{I}\mathbb{S}\mathbb{P}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Together with $\bm{A}\in\mathsf{K}\subseteq\mathbb{Q}(\mathsf{M})$, this yields $\bm{A}\in\mathbb{I}\mathbb{S}\mathbb{P}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Therefore, there exist $\{\bm{B}_{i}:i\in I\}\subseteq\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$ and an embedding $h\colon\bm{A}\to\prod_{i\in I}\bm{B}_{i}$. By Claim 8.12 there exists also $\{\bm{C}_{i}:i\in I\}\subseteq\mathsf{K}$ such that $\bm{B}_{i}\leqslant\bm{C}_{i}$ and $\langle p_{i}(h(a_{1})),\dots,p_{i}(h(a_{n}))\rangle\in\mathsf{dom}(f^{\bm{C}_{i}})$ for each $i\in I$.
@@
    if $\mathsf{K}$ is a quasivariety, then $\mathsf{ext}_{pp}(\mathsf{K})=\mathsf{ext}_{pp}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}},\mathsf{K})$.
@@
(i): Let $\mathsf{K}$ be a universal class and recall from 2.15 that $\mathsf{K}=\mathbb{U}(\mathsf{K}^{\textup{fg}})$. Therefore, $\mathsf{K}$ is elementary and $\mathsf{K}^{\textup{fg}}\subseteq\mathsf{K}\subseteq\mathbb{U}(\mathsf{K}^{\textup{fg}})$. Consequently, we can apply Proposition 8.11(i) to the case where $\mathsf{M}=\mathsf{K}^{\textup{fg}}$, obtaining $\mathsf{ext}(\mathsf{K})=\mathsf{ext}(\mathsf{K}^{\textup{fg}},\mathsf{K})$.
(ii): The inclusion from left to right is straightforward. To prove the other inclusion, observe that 2.16 guarantees that $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Moreover, $\mathbb{P}(\mathsf{K})\subseteq\mathsf{K}$ because $\mathsf{K}$ is a quasivariety. Therefore, we can apply 8.11(ii), obtaining $\mathsf{ext}_{pp}(\mathsf{K})=\mathsf{ext}_{pp}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}},\mathsf{K})$. ∎
@@
Observe that $\varphi$ is functional in $\mathsf{M}$ by assumption. As $\varphi$ is a pp formula, we can apply 3.11, obtaining that $\varphi$ is functional in $\mathbb{Q}(\mathsf{M})$ as well. Since $\mathsf{M}\subseteq\mathsf{K}$ and $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}\subseteq\mathbb{S}(\mathsf{M})$, we have $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ by Proposition 2.16. Hence, we conclude that $\varphi$ defines an implicit operation $f$ of $\mathsf{K}$ which, moreover is extendable by Theorem 8.13(ii). ∎
@@
Let $\mathsf{K}$ be a relatively filtral quasivariety. Then there exists a quaternary $f\in\mathsf{ext}_{eq}(\mathsf{K})$ such that $f^{\bm{A}}$ is total and coincides with the quaternary discriminator function on $A$ for each $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$.
@@
Therefore, for all $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$ and $a,b,c,d\in A$ there exists a unique $e\in A$ such that $\bm{A}\vDash\varphi(a,b,c,d,e)$. Then 8.14 implies that $\varphi$ defines an extendable implicit operation $f$ of $\mathsf{K}$. In view of the above display, $f^{\bm{A}}$ coincides with $d_{A}$ for each $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$. ∎
@@
In 3.16 we proved that there exists an implicit operation of the quasivariety $\mathsf{RCRing}$ of reduced commutative rings that coincides with the operation of “taking weak inverses” in fields. We now show that this operation is extendable.
@@
There exists a unary $f\in\mathsf{ext}_{eq}(\mathsf{RCRing})$ such that $f^{\bm{A}}$ is total and coincides with the operation of taking weak inverses for each field $\bm{A}$.
@@
Let $f$ be the implicit operation of $\mathsf{RCRing}$ given by 3.18. Then $f^{\bm{A}}$ is total and coincides with the operation of taking weak inverses for every field $\bm{A}$. Moreover, $f$ is defined by a conjunction of equations, whence $f\in\mathsf{imp}_{eq}(\mathsf{RCRing})$. Therefore, it suffices to prove that $f$ is extendable. To this end, recall from 3.17 that $\mathsf{RCRing}=\mathbb{Q}(\mathsf{Field})$, where $\mathsf{Field}$ is the class of fields. As $f^{\bm{A}}$ is total for each field $\bm{A}$, we can apply 8.11(ii) (taking $\mathsf{M}=\mathsf{Fields}$), obtaining that $f$ is extendable. ∎
@@
Recall from 3.19 that “taking relative complements” defines an implicit operation of the variety $\mathsf{DL}$ of distributive lattices and that “taking complements” defines an implicit operation of the variety $\mathsf{bDL}$ of bounded distributive lattices. We show that these operations are extendable.
@@
    the operation of taking relative complements in $\mathsf{DL}$ is a ternary member of $\mathsf{ext}_{eq}(\mathsf{DL})$;
@@
    the operation of taking complements in $\mathsf{bDL}$ is a unary member of $\mathsf{ext}_{eq}(\mathsf{bDL})$.
@@
We detail only the proof of (i), as the proof of (ii) is analogous. In view of 3.20, it suffices to show that the implicit operation $f$ of “taking relative complements” of $\mathsf{DL}$ is extendable. Let $\bm{D}_{2}$ be the two-element lattice and observe that $f^{\bm{D}_{2}}$ it total. As $\bm{D}_{2}$ is (up to isomorphism) the only subdirectly irreducible member of $\mathsf{DL}$ (see, e.g., [11, Ex. 3.19]), we obtain that $f^{\bm{A}}$ it total for each $\bm{A}\in\mathsf{DL}_{\textsc{si}}$. Therefore, we can apply Theorem 8.13(ii), obtaining that $f$ is extendable. ∎
@@
There exists a binary $f\in\mathsf{ext}_{eq}(\mathsf{Hilbert})$ such that $f^{\bm{A}_{\to}}$ is total and coincides with $\land^{\bm{A}}$ for each Heyting algebra $\bm{A}$.
@@
Let $\mathsf{M}$ be the class of implication reducts of Heyting algebras. Since the variety $\mathsf{Hilbert}$ is the class of implication subreducts of Heyting algebras, we have $\mathsf{Hilbert}=\mathbb{Q}(\mathsf{M})$. Moreover, observe that $\varphi$ is functional in $\mathsf{M}$ by Claim 8.23. As $\varphi$ is a conjunction of equations, we can apply 3.11, obtaining that $\varphi$ defines an implicit operation $f$ of $\mathsf{Hilbert}$. In view of Claim 8.23, it only remains to show that $f$ is extendable. To this end, consider a Hilbert algebra $\bm{A}$ and $a,b\in A$. Then $\bm{A}$ is a subreduct of a Heyting algebra $\bm{B}$, that is, $\bm{A}\leqslant\bm{B}_{\to}$. Since $f^{\bm{B}_{\to}}$ is total by Claim 8.23, we conclude that $f$ is extendable. ∎
@@
There exists a binary $f\in\mathsf{ext}_{eq}(\mathsf{PDL})$ such that $f^{\bm{A}}$ is total and coincides with $\to^{\bm{A}}$ for each finitely generated $\bm{A}\in\mathsf{PDL}_{\textsc{si}}$.

### v2

@@ 8. Extendable implicit operations
$\mathsf{ext}_{\textsc{pp}}(\mathsf{M},\mathsf{K})=\mathsf{ext}(\mathsf{M},\mathsf{K})\cap\mathsf{imp}_{\textsc{pp}}(\mathsf{K})\,\,\text{ and }\,\,\mathsf{ext}_{\textsc{eq}}(\mathsf{M},\mathsf{K})=\mathsf{ext}(\mathsf{M},\mathsf{K})\cap\mathsf{imp}_{\textsc{eq}}(\mathsf{K}).$
When $\mathsf{M}=\mathsf{K}$, we write $\mathsf{ext}(\mathsf{K}),\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$, and $\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ instead of $\mathsf{ext}(\mathsf{K},\mathsf{K}),\mathsf{ext}_{\textsc{pp}}(\mathsf{K},\mathsf{K})$, and $\mathsf{ext}_{\textsc{eq}}(\mathsf{K},\mathsf{K})$. Moreover, when an implicit operation is in $\mathsf{ext}(\mathsf{K})$, we simply say it is extendable.
@@
The proof of Theorem 8.4 hinges on the following observation.
@@
Recall from Theorem 3.9 that each implicit operation $f$ of $\mathsf{K}$ is defined by an existential positive formula $\varphi_{f}$. Then consider the following set of formulas in the language of $\mathsf{K}$ expanded with fresh constants $\{c_{a}:a\in A\}$ for the elements of $A$:
@@
Now, we proceed to prove the first part of the statement of Theorem 8.4. Consider $\bm{A}\in\mathsf{K}$. We will define a sequence $\{\bm{A}_{i}:i\in\mathbb{N}\}$ of members of $\mathsf{K}$. First, let $\bm{A}_{0}=\bm{A}$. Then suppose $\bm{A}_{i}\in\mathsf{K}$ has already been defined. By Claim 8.6 there exists $\bm{A}_{i+1}\in\mathsf{K}$ with $\bm{A}_{i}\leqslant\bm{A}_{i+1}$ and $A_{i}^{n}\subseteq\mathsf{dom}(f^{\bm{A}_{i+1}})$ for each $n$-ary $f\in\mathsf{ext}(\mathsf{K})$. By definition the sequence $\{\bm{A}_{i}:i\in\mathbb{N}\}$ constructed in this way is such that
@@
Now, as $\mathsf{K}$ is a universal class, it is closed under unions of chains of algebras by Proposition 2.4. Therefore, the union $\bm{B}$ of the chain in the above display belongs to $\mathsf{K}$. Furthermore, $\bm{A}=\bm{A}_{0}\leqslant\bm{B}$. To conclude the proof of the first part of the statement, it only remains to show that $f^{\bm{B}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. To this end, let $f\in\mathsf{ext}(\mathsf{K})$ be $n$-ary and $b_{1},\dots,b_{n}\in B$. As $\bm{B}$ is the union of the chain in the above display, there exists $i\in\mathbb{N}$ such that $b_{1},\dots,b_{n}\in A_{i}$. By the definition of $\bm{A}_{i+1}$ we have $\langle b_{1},\dots,b_{n}\rangle\in A_{i}^{n}\subseteq\mathsf{dom}(f^{\bm{A}_{i+1}})$. As $\bm{A}_{i+1}\leqslant\bm{B}$, we can apply Proposition 8.1, obtaining $\langle b_{1},\dots,b_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$. Hence, $f^{\bm{B}}$ is a total operation, as desired.
To prove the second part of the statement of Theorem 8.4, suppose that $\mathsf{K}$ is a quasivariety and consider $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$. In view of the first part of the statement of Theorem 8.4, there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $f^{\bm{B}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. By Proposition 8.5 there also exist $\bm{C}\in\mathsf{K}_{\textsc{rsi}}$ with $\bm{A}\leqslant\bm{C}$ and a surjective homomorphism $h\colon\bm{B}\to\bm{C}$. To conclude the proof, it only remains to show that $f^{\bm{C}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. To this end, consider an $n$-ary $f\in\mathsf{ext}(\mathsf{K})$ and $c_{1},\dots,c_{n}\in C$. Since $h\colon\bm{B}\to\bm{C}$ is surjective, there exist $b_{1},\dots,b_{n}\in B$ such that $h(b_{j})=c_{j}$ for each $j\leqslant n$. Recall that $f^{\bm{B}}$ is total because $f\in\mathsf{ext}(\mathsf{K}^{\textup{fg}},\mathsf{K})$. Therefore, $\langle b_{1},\dots,b_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$. As $f$ is an implicit operation and $h$ a homomorphism, we conclude that $\langle c_{1},\dots,c_{n}\rangle=\langle h(b_{1}),\dots,h(b_{n})\rangle\in\mathsf{dom}(f^{\bm{C}})$. ∎
The following is a consequence of Theorem 8.4.
@@
Given a universal class $\mathsf{K}$, the classes $\mathsf{ext}(\mathsf{K})$ and $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ are closed under composition.
@@
Consider an $n$-ary $g\in\mathsf{ext}(\mathsf{K})$ and $m$-ary $f_{1},\dots,f_{n}\in\mathsf{ext}(\mathsf{K})$. Let $\bm{A}\in\mathsf{K}$. By Theorem 8.4 there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $g^{\bm{B}},f_{1}^{\bm{B}},\dots,f_{n}^{\bm{B}}$ are total. It then follows from the definition of composition that $g(f_{1},\dots,_{n})^{\bm{B}}$ is also total. Therefore, $\mathsf{ext}(\mathsf{K})$ is closed under composition. As $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ is closed under composition by Proposition 3.13, we obtain that $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ is also closed under composition because $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})=\mathsf{ext}(\mathsf{K})\cap\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. ∎
@@
Recall from Example 4.16 that the class of cancellative commutative monoids forms a quasivariety, which we denote by $\mathsf{CCMon}$. The importance of cancellative commutative monoids is due to the following well-known result (see, e.g., [84, pp. 39–40]).
@@
    if $\mathbb{P}(\mathsf{K})\subseteq\mathsf{K}\subseteq\mathbb{Q}(\mathsf{M})$, then $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})=\mathsf{ext}_{\textsc{pp}}(\mathsf{M},\mathsf{K})$.
@@
(i): Suppose that $\mathsf{K}\subseteq\mathbb{U}(\mathsf{M})$. The inclusion from left to right follows from Remark 8.3. To prove the other inclusion consider $f\in\mathsf{ext}(\mathsf{M},\mathsf{K})$, $\bm{A}\in\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. By Theorem 2.2 we have $\mathbb{U}(\mathsf{M})=\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Together with $\bm{A}\in\mathsf{K}\subseteq\mathbb{U}(\mathsf{M})$, this yields $\bm{A}\in\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Therefore, there exist $\bm{B}\in\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$ and an embedding $h\colon\bm{A}\to\bm{B}$. By Claim 8.12 there exists also $\bm{C}\in\mathsf{K}$ such that $\bm{B}\leqslant\bm{C}$ and $\langle h(a_{1}),\dots,h(a_{n})\rangle\in\mathsf{dom}(f^{\bm{C}})$.
@@
(ii): Suppose that $\mathbb{P}(\mathsf{K})\subseteq\mathsf{K}\subseteq\mathbb{Q}(\mathsf{M})$. The inclusion from left to right follows from Remark 8.3. To prove the other inclusion consider $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{M},\mathsf{K})$, $\bm{A}\in\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. Assume that $f$ is defined by a pp formula $\varphi$. By Theorem 2.2 we have $\mathbb{Q}(\mathsf{M})=\mathbb{I}\mathbb{S}\mathbb{P}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Together with $\bm{A}\in\mathsf{K}\subseteq\mathbb{Q}(\mathsf{M})$, this yields $\bm{A}\in\mathbb{I}\mathbb{S}\mathbb{P}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Therefore, there exist $\{\bm{B}_{i}:i\in I\}\subseteq\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$ and an embedding $h\colon\bm{A}\to\prod_{i\in I}\bm{B}_{i}$. By Claim 8.12 there exists also $\{\bm{C}_{i}:i\in I\}\subseteq\mathsf{K}$ such that $\bm{B}_{i}\leqslant\bm{C}_{i}$ and $\langle p_{i}(h(a_{1})),\dots,p_{i}(h(a_{n}))\rangle\in\mathsf{dom}(f^{\bm{C}_{i}})$ for each $i\in I$.
@@
    if $\mathsf{K}$ is a quasivariety, then $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})=\mathsf{ext}_{\textsc{pp}}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}},\mathsf{K})$.
@@
(i): Let $\mathsf{K}$ be a universal class and recall from Proposition 2.15 that $\mathsf{K}=\mathbb{U}(\mathsf{K}^{\textup{fg}})$. Therefore, $\mathsf{K}$ is elementary and $\mathsf{K}^{\textup{fg}}\subseteq\mathsf{K}\subseteq\mathbb{U}(\mathsf{K}^{\textup{fg}})$. Consequently, we can apply Proposition 8.11(i) to the case where $\mathsf{M}=\mathsf{K}^{\textup{fg}}$, obtaining $\mathsf{ext}(\mathsf{K})=\mathsf{ext}(\mathsf{K}^{\textup{fg}},\mathsf{K})$.
(ii): The inclusion from left to right is straightforward. To prove the other inclusion, observe that Proposition 2.16 guarantees that $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}})$. Moreover, $\mathbb{P}(\mathsf{K})\subseteq\mathsf{K}$ because $\mathsf{K}$ is a quasivariety. Therefore, we can apply Proposition 8.11(ii), obtaining $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})=\mathsf{ext}_{\textsc{pp}}(\mathsf{K}_{\textsc{rsi}}^{\textup{fg}},\mathsf{K})$. ∎
@@
Observe that $\varphi$ is functional in $\mathsf{M}$ by assumption. As $\varphi$ is a pp formula, we can apply Corollary 3.11, obtaining that $\varphi$ is functional in $\mathbb{Q}(\mathsf{M})$ as well. Since $\mathsf{M}\subseteq\mathsf{K}$ and $\mathsf{K}_{\textsc{rsi}}^{\textup{fg}}\subseteq\mathbb{S}(\mathsf{M})$, we have $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ by Proposition 2.16. Hence, we conclude that $\varphi$ defines an implicit operation $f$ of $\mathsf{K}$ which, moreover is extendable by Theorem 8.13(ii). ∎
@@
Let $\mathsf{K}$ be a relatively filtral quasivariety. Then there exists a quaternary $f\in\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ such that $f^{\bm{A}}$ is total and coincides with the quaternary discriminator function on $A$ for each $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$.
@@
Therefore, for all $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$ and $a,b,c,d\in A$ there exists a unique $e\in A$ such that $\bm{A}\vDash\varphi(a,b,c,d,e)$. Then Corollary 8.14 implies that $\varphi$ defines an extendable implicit operation $f$ of $\mathsf{K}$. In view of the above display, $f^{\bm{A}}$ coincides with $d_{A}$ for each $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$. ∎
@@
In Example 3.16 we proved that there exists an implicit operation of the quasivariety $\mathsf{RCRing}$ of reduced commutative rings that coincides with the operation of “taking weak inverses” in fields. We now show that this operation is extendable.
@@
There exists a unary $f\in\mathsf{ext}_{\textsc{eq}}(\mathsf{RCRing})$ such that $f^{\bm{A}}$ is total and coincides with the operation of taking weak inverses for each field $\bm{A}$.
@@
Let $f$ be the implicit operation of $\mathsf{RCRing}$ given by Theorem 3.18. Then $f^{\bm{A}}$ is total and coincides with the operation of taking weak inverses for every field $\bm{A}$. Moreover, $f$ is defined by a conjunction of equations, whence $f\in\mathsf{imp}_{\textsc{eq}}(\mathsf{RCRing})$. Therefore, it suffices to prove that $f$ is extendable. To this end, recall from Theorem 3.17 that $\mathsf{RCRing}=\mathbb{Q}(\mathsf{Field})$, where $\mathsf{Field}$ is the class of fields. As $f^{\bm{A}}$ is total for each field $\bm{A}$, we can apply Proposition 8.11(ii) (taking $\mathsf{M}=\mathsf{Fields}$), obtaining that $f$ is extendable. ∎
@@
Recall from Example 3.19 that “taking relative complements” defines an implicit operation of the variety $\mathsf{DL}$ of distributive lattices and that “taking complements” defines an implicit operation of the variety $\mathsf{bDL}$ of bounded distributive lattices. We show that these operations are extendable.
@@
    the operation of taking relative complements in $\mathsf{DL}$ is a ternary member of $\mathsf{ext}_{\textsc{eq}}(\mathsf{DL})$;
@@
    the operation of taking complements in $\mathsf{bDL}$ is a unary member of $\mathsf{ext}_{\textsc{eq}}(\mathsf{bDL})$.
@@
We detail only the proof of (i), as the proof of (ii) is analogous. In view of Theorem 3.20, it suffices to show that the implicit operation $f$ of “taking relative complements” of $\mathsf{DL}$ is extendable. Let $\bm{D}_{2}$ be the two-element lattice and observe that $f^{\bm{D}_{2}}$ it total. As $\bm{D}_{2}$ is (up to isomorphism) the only subdirectly irreducible member of $\mathsf{DL}$ (see, e.g., [11, Ex. 3.19]), we obtain that $f^{\bm{A}}$ it total for each $\bm{A}\in\mathsf{DL}_{\textsc{si}}$. Therefore, we can apply Theorem 8.13(ii), obtaining that $f$ is extendable. ∎
@@
There exists a binary $f\in\mathsf{ext}_{\textsc{eq}}(\mathsf{Hilbert})$ such that $f^{\bm{A}_{\to}}$ is total and coincides with $\land^{\bm{A}}$ for each Heyting algebra $\bm{A}$.
@@
Let $\mathsf{M}$ be the class of implication reducts of Heyting algebras. Since the variety $\mathsf{Hilbert}$ is the class of implication subreducts of Heyting algebras, we have $\mathsf{Hilbert}=\mathbb{Q}(\mathsf{M})$. Moreover, observe that $\varphi$ is functional in $\mathsf{M}$ by Claim 8.23. As $\varphi$ is a conjunction of equations, we can apply Corollary 3.11, obtaining that $\varphi$ defines an implicit operation $f$ of $\mathsf{Hilbert}$. In view of Claim 8.23, it only remains to show that $f$ is extendable. To this end, consider a Hilbert algebra $\bm{A}$ and $a,b\in A$. Then $\bm{A}$ is a subreduct of a Heyting algebra $\bm{B}$, that is, $\bm{A}\leqslant\bm{B}_{\to}$. Since $f^{\bm{B}_{\to}}$ is total by Claim 8.23, we conclude that $f$ is extendable. ∎
@@
There exists a binary $f\in\mathsf{ext}_{\textsc{eq}}(\mathsf{PDL})$ such that $f^{\bm{A}}$ is total and coincides with $\to^{\bm{A}}$ for each finitely generated $\bm{A}\in\mathsf{PDL}_{\textsc{si}}$.